OpenStudy (anonymous):

determine the dimensions of a rectangular solid (with a square base) with maximun volume if its surface area is 337.5 square centimeters.

OpenStudy (cwrw238):

Volume V = a^2h where a = length of a side of the base and h = height surface area 2a^2 + 4ah = 337.5 from this h = (337.5 - 2a^2) / 4a therefore the volume V = a^2(337. 5 - 2a^2) / 4a V = 0.25a(337.5 - 2a^2) V = 84.375 - 0.5a^3

OpenStudy (cwrw238):

now find dV/da and equate to 0 solve for a and then find h validate that this is a maximum

OpenStudy (cwrw238):

oops sorry theres an error in the last line correct is V = 84.375a - 0.5a^3